lifewmylilweeb:

@Intrinix One x-intercept for a parabola is at the point (-3,0). Use the factor method to find the other x-intercept for the parabola defined by this equation: y=-x^2-5x-6 Seperate the values with a comma.

lifewmylilweeb:

can you show me how to the x-intercept now please

axie:

@extrinix

lifewmylilweeb:

i have no idea do i use the quadrativ formula for this one

lifewmylilweeb:

what do i do next because i dont remember doing it this way in class

lifewmylilweeb:

i think they made us solve it using the quadratic formula- [y=\frac{ -b+\sqrt{b^2-4(ac)} }{ 2(a) }\]

Extrinix:

\(y=\frac{-b+\sqrt{b^2-4(ac)}}{2a}\) \(y=\frac{-(-5)+\sqrt{(-5)^2-4(-1)(-6)}}{2(-1)}\) So we would simplify it \(y=\frac{5+\sqrt{1}}{-2}\) What would this equal?

Extrinix:

See how that wouldn't work? Because that's looking for the \(y-int\), not \(x-int\)

lifewmylilweeb:

it would equal -3

Extrinix:

Exactly, and that's the y-int we got the other day

lifewmylilweeb:

oh

Extrinix:

So we would do it by factoring \(0=-x^2-5x-6\) Add \(-x^2-5x-6\) to both sides (to flip it) \(x^2+5x+6=0\) So what multiplies to get 6 and adds to get 5? __ \(\times\) __ = 6 __ \(+\) __ = 5

lifewmylilweeb:

3*2=6 3+2=5

Extrinix:

Which would give us \((x+3)(x+2)\) So now we set each \(()\) equal to zero and solve for both of them \(x+3=0\) | \(x+2=0\) So what would you get?

lifewmylilweeb:

-3,-2

Extrinix:

So those are your two x-intercepts

lifewmylilweeb:

so (-3,-2)

Extrinix:

And you already have \((-3,0)\), so what would the other one be?

Extrinix:

No, x-int is \((x,0)\), remember?

lifewmylilweeb:

ok (-2,0

Extrinix:

Correct